Determine the reactions at the bearing supports A, B, and C of the shaft, then draw the shear and moment diagrams. EI is constant. Each bearing exerts only vertical reactions on the shaft.
400 N 400 N
Answer #1

Free Body Diagram HOON 400 N 1 lm Im ↑ 1m Im f Ay By Cy Take By as redundant. And wing method of super position, find deflectThe compatibility condition is not deflection at B=0. .: SBI + SB2 +8B3 = 0 = -366.61 - 366.64 + 1.333 By =0 GL EI ст . By =RB is taken as redundant and using compatibility equation and method of super position, the reactions are found:

RA= 125N

RB= 550N

RC= 125N